what is 2/5 to the third power
Answer by 2897696(96)

we know that x^2=x*x, y^7=y*y*y*y*y*y*y,or (y/u)^3=(y/u)*(y/u)*(y/u), so
(2/5)^3=(2/5)*(2/5)*(2/5).
now we multiply the numerator and the denominator together.
(2*2*2)/(5*5*5)=8/125
so, 2/5 to the third power is 8/125